OpenStudy (rissyroo118):

WILL MEDAL & FAN ? :) If a die is rolled twice, what is the probability of rolling a 5 and then a 2? A)1/36 B)1/3 C)1/6 D)2/36

OpenStudy (anonymous):

what is the probability of rolling a five?

OpenStudy (anonymous):

\[\frac{ 1 }{ 6 }\times \frac{ 1 }{ 6 }=?\]

OpenStudy (rissyroo118):

1/36